in the equation p(x)=12x-100 what is the meaning of the number 12 in the formula?
what is the meaning of -100 in the formula?
find the profit for 5 different vaules of x?

Answer by stanbon(75887)

in the equation p(x)=12x-100
what is the meaning of the number 12 in the formula?
Ans: It is the slope of the equation.
It is the amount p increases when x increases by 1.
-------------------------
what is the meaning of -100 in the formula?
It is the p-intercept.
It is the value of P when x = 0
-------
find the profit for 5 different vaules of x?
f(1) = 12-100 = -88
f(5) = 12*5-100 = -30
etc.
Cheers,
Stan H.
============